OpenStudy (anonymous):

A spring with a spring constant of 100 pounds per foot is loaded with 1-pound weight and brought to equilibrium. It is then stretched an additional 1 inch and released. Find the equation of motion, the amplitude, and the period. Neglect friction.

OpenStudy (anonymous):

before we start, may I know if the spring was placed vertically or horizontally?

OpenStudy (anonymous):

^ that should be irrelevant

OpenStudy (anonymous):

No derivation? lol Could have just given the \(T=2\pi \sqrt{\frac{k}{m}}\) and get everything else from this.

OpenStudy (anonymous):

What? I just meant that everything is the same either way.

OpenStudy (anonymous):

oh. i see. lol